Find the probability that the sum is as stated when a pair of dice is rolled. (Enter your answers as fractions.)
(a) odd and doubles
(b) odd or doubles

Found 2 solutions by Edwin McCravy, mananth:
Answer by Edwin McCravy(20054)

   
Here are all 36 possible rolls:

(1,1) (2,1) (3,1) (4,1) (5,1) (6,1)
(1,2) (2,2) (3,2) (4,2) (5,2) (6,2)
(1,3) (2,3) (3,3) (4,3) (5,3) (6,3)
(1,4) (2,4) (3,4) (4,4) (5,4) (6,4)
(1,5) (2,5) (3,5) (4,5) (5,5) (6,5)
(1,6) (2,6) (3,6) (4,6) (5,6) (6,6)
(a) odd and doubles
Count how many BOTH have odd sums AND ALSO are doubles. Make a fraction with
that number as the numerator and 36 as the denominator. If it will reduce, do
so.
(b) odd or doubles
Count how many EITHER have odd sums OR are doubles. Make a fraction with that 
number as the numerator and 36 as the denominator. If it will reduce, do so.

Edwin

Answer by mananth(16946)

Answer on Question #66364 - Math - Statistics and ProbabilityQuestionFind the probability that the sum is as stated when a pair of dice is rolled. (Enter your answers as fractions.)(a)odd and doubles(b)odd or doubles
a) Sample space = Total number of outcomes = 36
Event of getting sum odds
A= (1,2),(1,4),(1,6),(2,1),(2,3),(2,5),(3,2).(3,4),
(3,6),(4,1),(4,3),(4,5),(5,2),(5,4),(5,6)(6,1),(6,3),(6,5)
= 18
Doubles =(1,1)(2,2)(3,3),(4,4),(5,5),(6,6)
=(6)
P (odd and double) = (0)
b) P(odd or doub;es ) = (18+6)/36 = 24/36= 2/3